OpenStudy (anonymous):

solve sin^2x+2sinx+1=0

OpenStudy (anonymous):

\[\sin^2x + 2 sinx + 1 = (sinx+1)^2\] \[(sinx+1)^2 = 0\] \[sinx + 1 = 0\] \[sinx = -1\] \[x = \arcsin(-1)\]

OpenStudy (anonymous):

x = 3pi/2 + 2 pi k where k is any integer (if i didn't forget my angles lol)

OpenStudy (anonymous):

different question.

OpenStudy (anonymous):

answer: http://adf.ly/85KnH